Given the trinomial square x^2-8x=4 solve for the solutions of x using the completing square method. Show all work. Thank you in advance for your help
Answer by stanbon(75887)

Given the trinomial square x^2-8x=4 solve for the solutions of x using the completing square method.
------
x^2-8x = 4
-----
x^2 -8x + 16 = 4+16
-----
Factor:
(x-4)^2 = 20
--------
x-4 = 2sqrt(5) or x-4 = -2sqrt(5)
--------
x = 4+2sqrt(5) or x = 4-2sqrt(5)
====================================
Cheers,
Stan H.